Can you please help,Von is thrice as old as his brother Jon. Four years ago, he was 4 years less than 5 times Jon�s age at that time. How old are they now?

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

v = von's age
j = jon's age.
v = 3j is your first equation.
v-4 = 5(j-4)-4 is your second equation.
you have 2 equations that need to be solved simultaneously.
they are:
v = 3j
v-4 = 5(j-4)-4
simplify the second equation to get:
v-4 = 5j-20-4 which simplifies further to:
v-4 = 5j-24
add 4 to both sides of this equation to get:
v = 5j-20
your 2 equations are now:
v = 3j
v = 5j-20
since both expressions on the right side of each equation are equal to v, you can set these expressions equal to each other to get:
3j = 5j-20
subtract 5j from both sides of this equation to get:
-2s = -20
divide both sides of this equation by -2 to get:
j = 10
your first equation was:
v = 3j
this makes v = 30
your second equation was:
v-4 = 5(j-4)-4
replace v with 30 and j with 10 to get:
30-4 = 5(10-4)-4 which simplifies to:
26 = 50-20-4 which simplifies further to:
26 = 30-4 which simplifies further to:
26 = 26
the values of 30 for v and 10 for j are good.

Answer by ikleyn(53751)

.
Can you please help,
Von is thrice as old as his brother Jon. Four years ago, he was 4 years less than 5 times Jon's age at that time.
How old are they now?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        This problem can be solved in hard way,  using two equations and writing many words,  as tutor  @Theo did in his post.

        It also can be solved in easy way,  using one equation and writing only the necessary words,  as I present it below.


Let x be the Jon age.
Then the Von age is 3x, according to the problem.


Four years ago, the Jon's age was x-4.
Four years ago, the Von's age was (3x-4).


From the problem, we have this equation

    3x-4 = 5(x-4) - 4.    (Four years ago, Von was 4 years less than 5 times Jon's age at that time).


It is an equation to solve. Simplify and find x

    3x - 4 = 5x - 20 - 4.


Collect the terms with 'x' on the right side; collect the constant terms on the left side

    -4 + 20 + 4 = 5x - 3x,

         20     =    2x

          x     =    20/2 = 10.


Thus, Jon is 10 years old;  Von is 3*10 = 30 years old.    <<<---===  ANSWER

Solved.   The solution is clear,  transparent and straightforward,  with no excessive words that only distract attention.